are the following pairs of lines parallel 2x - 4y = 16 and 3x = 6y + 18
Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(9144)

-2/-4=1/2
-3/-6=1/2
yes

Answer by stanbon(57967)

are the following pairs of lines parallel
2x - 4y = 16
y = (1/2)x -4
-------------------
3x = 6y + 18
y = (1/2)x-3
-----
Slopes are the same and intercepts are different, so the lines ae parallel.
-----------------
Cheers,
Stan H.
-----------------